# HLStrans: Dataset for LLM-Driven C-to-HLS Hardware Code Synthesis

### v1

@@ A.1 HLS Code Transformations
In our datasets, we apply a suite of code-reconstruction techniques designed to optimize memory access patterns, alleviate computational bottlenecks, and resolve loop dependencies. By restructuring data flow and exploiting hardware parallelism, these methods boost throughput and shorten overall execution time. Table 6 list the main Code Restructuring adopted in our dataset.
@@
Our dataset features an extensive catalog of HLS pragmas ranging from memory-access directives (array partitioning, streaming) through loop-level transformations (unrolling, merging, tiling) to fine-grained pipeline controls (initiation interval tuning, dataflow regions). By systematically applying and combining these pragmas, these directives empower automated HLS flows to tailor synthesized hardware for domain-specific latency, throughput, and area requirements making our dataset a valuable reference for exploring pragma-driven performance tuning. Table 7 introduce these applied pragma optimizations.
@@
Memory type Specify the on-chip storage type (BRAM/URAM/SMALL_RAM) #pragma HLS RESOURCE variable=buf core=RAM_2P
@@
Dataflow Enable task-level parallelism between functions #pragma HLS DATAFLOW
@@
struct/union struct { ap_uint<…> field; … } with bitfields Precisely specify field bit-widths and alignment, eliminate padding
pointer/array hls::stream<T> Map to hardware FIFO streams for stream transmission
@@ A.1.4 HLS Code Optimization: Transformation of Functions
std::exp(x) hls::exp(x) Synthesizes an exponential function hardware block (LUT-based)
std::log(x) hls::log(x) Provides a hardware-friendly implementation of natural logarithm
@@
High-level synthesis (HLS) cannot synthesize all idiomatic C constructs directly. To enable hardware generation, we must refactor unsupported patterns like dynamic memory allocation, recursion, and pointer arithmetic into HLS-compliant coding styles that the tool can analyze and map to on-chip resources. Table 10 lists these common transformations.
@@
A.2 Dataset generations
A.2.1 Dataset collections
We began by collecting a comprehensive set of human-optimized HLS implementations from GitHub repositories and real-world FPGA applications. Our curated corpus spans diverse application domains and covers all five transformation categories detailed in Appendix A.1. Figure 5 visualizes the distribution of programs across these five tasks, and the source kernels themselves fall into seven distinct application categories. This rich, well-balanced dataset provides the broad coverage of real-world HLS patterns necessary to train and evaluate LLM’s hardware synthesis ability.
Figure 5: dataset program source distribution
A.2.2 Dataset Augmentation
Retrieval-Augmented Generation: To broaden the range of HLS code-transformation techniques that our LLM can learn, we built an automated framework (see Appendix A.2) that programmatically generates optimized variants via Monte Carlo Tree Search. Central to this system is a Retrieval-Augmented Generation (RAG) table of optimization strategies including code-reconstruction patterns, directive (pragma) insertions, data-type adaptations, and function-level transformations each entry pairing a concise description with a few-shot example illustrating the baseline code and its optimized counterpart listed in A.1. During search, these RAG-driven actions guide the MCTS policy to apply specific transformations, yielding a diverse corpus of HLS kernels ready for LLM fine-tuning and evaluation. One kind of Retrieval-Augmented strategies is shown in Fig.6 and prompt template is shown in Fig.7.
Figure 6: Example of Retrieval-Augmented strategies in MCTS framework
@@ A.2.2 Dataset Augmentation
Framework Evaluation: We evaluated our dataset-augmentation framework on the widely adopted Rodinia benchmark suite [38], using a Xilinx Alveo U55C FPGA board running at a 300 MHz kernel clock. Our goal was to measure how effectively our MCTS-based sampler could guide Deepseek-R1 and GPT-4o toward highly optimized HLS kernels.
Table 11 summarizes the Best@1 kernel runtimes (in milliseconds) across twelve diverse applications, comparing four configurations: Baseline (The unmodified, compiler-generated HLS implementation), HLSPilot [7] (A recent LLM based optimization framework), GPT-4o with our framework and Deepseek-R1 with our framework. Our results reveal several key findings:
  • Consistent Improvement over previous work. In every benchmark, both of our enhanced pipelines outperform HLSPilot, demonstrating that the combination of large-model code generators with MCTS exploration yields more hardware-efficient HLS designs.
@@
  • Robust Gains Across Diverse Kernels. From compute-bound codes such as kmeans, mgvf, and streamcluster, to memory-sensitive workloads like hotspot and nw, our framework consistently identifies and applies scheduling, pipelining, and data-partitioning transformations that exploit the parallelism and memory hierarchy of the Alveo U55C.
@@
Application Baseline HLSPilot [7] Ours
@@ A.2.3 Dataset Examples
This section presents three real HLS code transformation pair examples: performance optimization (Code restructuring and Directive insertion), synthesizability correction (Code repair), and adaptation from C-style to HLS-style code (Data-type adaptation and transformation of functions). The dataset is hosted at https://huggingface.co/datasets/qingyun777yes/HLStrans.
@@ 1. Performance Optimization
Figures 8(a) and 8(b) show a simple K-Nearest Neighbors (KNN) implementation before and after HLS optimization. The optimized version achieves better performance due to improved pipelining, parallelism, and memory optimization.
@@ 2. Synthesizability Transformation
Figures 9(a) and 9(b) illustrate the transformation from a non-synthesizable function into a valid HLS-compatible version.
@@ 3. C-style to HLS-style Conversion
Figures 10(a) and 10(b) demonstrate how C-style data types and functions can be adapted into HLS-friendly forms.
@@ A.4.1 HLS Synthesis Results Example
While HLS synthesis cannot yield perfectly accurate timing or resource-utilization numbers, it provides essential estimates for comparing design variants. Figures 14(a) and 14(b) show the synthesis reports for the unoptimized and optimized KNN kernels, respectively, targeting a Xilinx Alveo U55C accelerator at a 300 MHz kernel clock.
The optimized design trades increased resource usage more DSP slices, flip-flops (FFs), and lookup tables (LUTs) for a dramatic fourfold reduction in latency cycles (from 2,097,324 cycles down to 508,479 cycles). In a real FPGA deployment, this corresponds to an end-to-end runtime of approximately 3.2 ms versus roughly 17 ms for the unoptimized kernel, while still remaining within resource budgets. We report acceleration based on the estimated latency cycles from the synthesis reports. However, HLS synthesis itself can be time-consuming, particularly for large or highly optimized designs.
@@ A.4.2 Finetune Results
With our real-world corpus, we reserve two C programs for the repair task and another 15 programs for the optimization task. The remaining 120 programs are split into training and validation sets. Figure 15(a) presents the validation loss curve for Qwen2.5-Coder-3B-Instruct, while Figure 15(b) shows the corresponding curve for Qwen2.5-Coder-7B-Instruct during fine-tuning. In both cases, the steadily decreasing loss demonstrates that fine-tuning effectively adapts the models to our dataset.

### v3

@@ A.1 HLS Code Transformations
In our datasets, we apply a suite of code‑reconstruction techniques designed to optimize memory access patterns, alleviate computational bottlenecks, and resolve loop dependencies. By restructuring data flow and exploiting hardware parallelism, these methods boost throughput and shorten overall execution time. Table 4 list the main Code Restructuring adopted in our dataset.
@@
Our dataset features an extensive catalog of HLS pragmas ranging from memory‑access directives (array partitioning, streaming) through loop‑level transformations (unrolling, merging, tiling) to fine‑grained pipeline controls (initiation interval tuning, dataflow regions). By systematically applying and combining these pragmas, these directives empower automated HLS flows to tailor synthesized hardware for domain‑specific latency, throughput, and area requirements making our dataset a valuable reference for exploring pragma‑driven performance tuning. Table 5 introduce these applied pragma optimizations.
@@
Memory type Specify the on‑chip storage type (BRAM/URAM/SMALL_RAM) #pragma HLS RESOURCE variable=buf core=RAM_2P
@@
Dataflow Enable task‑level parallelism between functions #pragma HLS DATAFLOW
@@
struct/union struct { ap_uint<…> field;} with bitfields Precisely specify field bit-widths and alignment, eliminate padding
pointer/array hls::stream<T> Map to hardware FIFO streams for streaming transmission
@@ A.1.4 HLS Code Optimization: Transformation of Functions
std::exp(x) hls::exp(x) Synthesizes an exponential function hardware block (LUT‑based)
std::log(x) hls::log(x) Provides a hardware‑friendly implementation of natural logarithm
@@
High‑level synthesis (HLS) cannot synthesize all idiomatic C constructs directly. To enable hardware generation, we must refactor unsupported patterns like dynamic memory allocation, recursion, and pointer arithmetic into HLS‑compliant coding styles that the tool can analyze and map to on-chip resources. Table 8 lists these common transformations.
@@
A.2 Dataset Augmentation
A.2.1 MCTS framework
Agent G: We propose to use an LLM as the agent that leverages its pretrained knowledge of hardware design and in‐context learning abilities.
Action A: At each time step $t$, the agent selects an action $a_{t}$, which corresponds to a prompt or transformation applied to the current HLS code. We define two complementary action types: RAG‑based actions retrieve optimization policies directly from our pre‐built table and accompanying code examples shown in Figure 6, leveraging retrieval‑augmented generation to surface proven transformations rapidly and reliably. Reasoning‑based actions with ReAct prompt (Yao et al., 2023), in contrast, analyze compiler warnings such as pipeline‑interval breaches or loop‑unroll violations and apply targeted code reforms by interpreting warning semantics within the current code context.
(a) Loop tiling Code examples by RAG
(b) LLM reasoning about environment warning and tool hint
Figure 6: Actions design of MCTS
Reward R: Rule‐based reward shaping has proven effective in guiding agent behavior in previous work Guo et al. (2025). In our framework, the reward function $R(s_{t},a_{t})$ is computed by applying rule‐based scoring to verification results and feedback provided by the HLS tool. A penalty of 2-2 is applied when the verification fails, and 1-1 if synthesis fails or the design exceeds resource constraints. A neutral reward of 0 is given when the transformed design performs worse than the original, while a reward of $1$ is granted when it performs better. If the design not only surpasses the original but also meets timing constraints, a higher reward of $2$ is assigned.
In our cases, the MCTS begins at the initial state $S_{0}$ (the root node), which is the naive HLS code. From a state $S_{t}$, the agent applies an optimization policy $\pi$, i.e., an action $a_{t}\in\mathcal{A}$, transitioning to the subsequent state $S_{t+1}$. MCTS consists of four key phases: selection, expansion, simulation, and backpropagation. $N(S_{t})$, the total number of times $S_{t}$ has been visited.
Selection: We employ the upper confidence bounds for trees (UCT) (Gelly & Wang, 2006) algorithm to choose nodes.
Expansion: From the current state $s_{t}$, generate one or more child nodes to explore untried actions.
Retrieval‐Augmented Generation: To broaden the range of HLS code‐transformation techniques that our LLM can learn, we built an automated framework (see Appendix A.2) that programmatically generates optimized variants via Monte Carlo Tree Search. Central to this system is a Retrieval‐Augmented Generation (RAG) table of optimization strategies including code‑reconstruction patterns, directive (pragma) insertions, data‑type adaptations, and function‑level transformations each entry pairing a concise description with a few‑shot example illustrating the baseline code and its optimized counterpart listed in A.1. During search, these RAG‑driven actions guide the MCTS policy to apply specific transformations, yielding a diverse corpus of HLS kernels ready for LLM fine‑tuning and evaluation. One kind of Retrieval‐Augmented strategies is shown in Fig.7 and prompt template is shown in Fig.9.
Figure 7: Example of Retrieval‐Augmented strategies in MCTS framework
@@ A.2.1 MCTS framework
Framework Evaluation: We evaluated our dataset‐augmentation framework on the widely adopted Rodinia benchmark suite (Che et al., 2009), using a Xilinx Alveo U55C FPGA board running at a 300 MHz kernel clock. Our goal was to measure how effectively our MCTS‐based sampler could guide Deepseek‐R1 and GPT-4o toward highly optimized HLS kernels.
Fig.7 shows the average success rate in the benchmarks. As the figure shows, their success rate is Qwen32B > Deepseek-R1 > GPT-4o > Qwen7B while Deepseek-R1 can achieve highest average speedup. The results show performance-increase will degrade the ability of LLM to produce the correct HLS code.
Figure 9: Success rate on different benchmarks with different models.
Table 9 summarizes the Best@1 kernel runtimes (in milliseconds) across twelve diverse applications, comparing four configurations: Baseline (The unmodified, compiler‐generated HLS implementation), HLSPilot (Xiong et al., 2024) (A recent LLM based optimization framework), GPT-4o with our framework and Deepseek-R1 with our framework. Our results reveal several key findings:
  • Consistent Improvement over previous work. In every benchmark, both of our enhanced pipelines outperform HLSPilot, demonstrating that the combination of large‐model code generators with MCTS exploration yields more hardware‐efficient HLS designs.
@@
  • Robust Gains Across Diverse Kernels. From compute‐bound codes such as kmeans, mgvf, and streamcluster, to memory‐sensitive workloads like hotspot and nw, our framework consistently identifies and applies scheduling, pipelining, and data‐partitioning transformations that exploit the parallelism and memory hierarchy of the Alveo U55C.
@@
Application Baseline HLSPilot Xiong et al. (2024) Ours
@@
A.2.2 Design Space Exploration.
The tool is an automated HLS design-space explorer that employs a genetic-algorithm optimizer to discover effective directive combinations—specifically loop pipelining, loop unrolling, and array partitioning—that maximize performance and resource efficiency. We traverse the search space with the NSGA-II algorithm (Deb et al., 2002) as implemented in the PyMOO library (Blank & Deb, 2020), chosen for its ability to escape local optima and rapidly converge to high-quality solutions. NSGA-II is executed for 24 generations with a population size of 40. Each generation performs three steps: (1) generate or initialize the population, (2) apply each candidate configuration to the source code using compiler B2 and synthesize with Xilinx Vitis, and (3) return the synthesis metrics to NSGA-II. Configurations that exceed device resources or demand prohibitive HLS runtimes (e.g., $\gtrsim 1$ hour) are deemed infeasible and discarded. Genetic operators are configured as follows: random sampling/selection (mutation sampling probability =0.1=0.1), simulated binary crossover (probability =1.0=1.0, $\eta=15$), and polynomial mutation ($\eta=20$); all other operator parameters use PyMOO defaults.
@@ A.2.3 Dataset Examples
This section presents three real HLS code transformation pair examples: performance optimization (Code restructuring and Directive insertion), synthesizability correction (Code repair), and adaptation from C-style to HLS-style code (Data-type adaptation and transformation of functions).
@@ 1. Performance Optimization
Figures 10(a) and 10(b) show a simple K-Nearest Neighbors (KNN) implementation before and after HLS optimization. The optimized version achieves better performance due to improved pipelining, parallelism, and memory optimization.
@@ 2. Synthesizability Transformation
Figures 11(a) and 11(b) illustrate the transformation from a non-synthesizable function into a valid HLS-compatible version.
@@ 3. C-style to HLS-style Conversion
Figures 12(a) and 12(b) demonstrate how C-style data types and functions can be adapted into HLS-friendly forms.
@@ A.4.1 HLS Synthesis Results Example
While HLS synthesis cannot yield perfectly accurate timing or resource‐utilization numbers, it provides essential estimates for comparing design variants. Figures 16(a) and 16(b) show the synthesis reports for the unoptimized and optimized KNN kernels, respectively, targeting a Xilinx Alveo U55C accelerator at a 300 MHz kernel clock.
The optimized design trades increased resource usage more DSP slices, flip-flops (FFs), and lookup tables (LUTs) for a dramatic fourfold reduction in latency cycles (from 2,097,324 cycles down to 508,479 cycles). In a real FPGA deployment, this corresponds to an end‐to‐end runtime of approximately 3.2 ms versus roughly 17 ms for the unoptimized kernel, while still remaining within resource budgets. We report acceleration based on the estimated latency cycles from the synthesis reports. However, HLS synthesis itself can be time-consuming, particularly for large or highly optimized designs.